OpenStudy (winsteria):

SOMEONE PLEASE HELP IF YOU DO I'LL GIVE YOU MEDAL AND FAN please help!!! Which related equation can be used to solve the equation? Column A 1. –8.7 = k + 6 2. –6 = k + 8.7 3. 8.7 = k + (–6) 4. 6 = k + 8.7 Column B A. k = –8.7 – 6 B. k = 6 – 8.7 C. k = 8.7 – (–6)

OpenStudy (winsteria):

you need to match up the column a answers with the column b answers

OpenStudy (winsteria):

does anyone know? :(

OpenStudy (winsteria):

@cody7744

OpenStudy (anonymous):

@Cydney_Morgan

OpenStudy (winsteria):

I wonder if anyone knows because im really confused :(

OpenStudy (anonymous):

it looks very confusing what grade are you in

OpenStudy (winsteria):

7th

OpenStudy (winsteria):

DOES ANYONE KNOW OMG :(

OpenStudy (triciaal):

for each question 1-4 doing them one at a time solve for k make k the subject remember to keep the balance what you do to one side of an equation also do to the other . you have 1 -4 k = whatever now look in column be and find the choice that matches. for example 2. –6 = k + 8.7 k is on the right to get k only -8.7 keep the balance so -8.7 from the left also you get -6 -8.7 = k which matches option A solution 2 A

OpenStudy (winsteria):

I still don't get how in the world i'm supposed to get the answers

OpenStudy (winsteria):

@GreeneBean @babyblue123

OpenStudy (winsteria):

does anyone know.

OpenStudy (anonymous):